OpenStudy (anonymous):

Can you help me solve this problem? 700(1000) + 1200R=1,000,000

OpenStudy (austinl):

Are we solving for "R"?

OpenStudy (anonymous):

Yerp

OpenStudy (austinl):

Ok, what do you think you should do first?

OpenStudy (anonymous):

Wait I think I know how to do it haha

OpenStudy (austinl):

If you get an answer, you can check it against mine.

OpenStudy (anonymous):

Would R be 250?

OpenStudy (austinl):

Close, but not quite.

OpenStudy (anonymous):

Okay so where do I start with solving this problem?

OpenStudy (austinl):

You need to multiply the 700 by the 1000, what do you get?

OpenStudy (anonymous):

I did that and got 700,000

OpenStudy (austinl):

Now, subtract that from the right side.

OpenStudy (anonymous):

I did that when I did it the first time and got 300,000

OpenStudy (austinl):

Oh my goodness, you were correct, I accidentally had an extra zero in one of my steps D: You divide that by 1200 and get 250. My bad :(

OpenStudy (anonymous):

Haha its cool

OpenStudy (anonymous):

Thanks for helping meh

OpenStudy (anonymous):

I have a few more I need help working through if you wanna help me with them

OpenStudy (anonymous):

:)

OpenStudy (austinl):

Sure, why not :)

OpenStudy (anonymous):

Do you mind if I just put them in the comments right here?

OpenStudy (austinl):

Nope, feel free :)

OpenStudy (anonymous):

Its not letting me paste the equation into the comments so Im going to attach a file to it if that's cool with you

OpenStudy (austinl):

Sure.

OpenStudy (anonymous):

There you go

OpenStudy (austinl):

Easier way yes @robtobey but we already found then answer. :) Even through my blunders :P

OpenStudy (anonymous):

^ XD

OpenStudy (austinl):

What am I looking at here?

OpenStudy (anonymous):

3 double equation word problems

OpenStudy (anonymous):

I don't know where to begin with these problems

OpenStudy (anonymous):

Its Algebra 1 by the way

OpenStudy (austinl):

I would say that you need to set P=62.10 and solve for x.

OpenStudy (anonymous):

Mkay give me a minute

OpenStudy (anonymous):

Okay Im trying to do it and Im getting lost haha can we go through it step by step on what I have to do?

OpenStudy (austinl):

Yep :)

OpenStudy (anonymous):

Thx <3

OpenStudy (anonymous):

Youre the best

OpenStudy (austinl):

We have, \(P=(20+0.5x)+0.15(20+0.5x)\) The most we can spend is 62.10, so we set P equal to this, \(62.10=(20+0.5x)+0.15(20+0.5x)\) Follow so far?

OpenStudy (anonymous):

Yep

OpenStudy (austinl):

Ok, we would want to distribute on the right side. \(0.15(20+0.5x)=3+.075x\) So now we have, \(62.10=20+0.5x+3+.075x\) Still follow?

OpenStudy (anonymous):

Im following

OpenStudy (anonymous):

Would we then combine like terms?

OpenStudy (austinl):

Yes we would! :D

OpenStudy (anonymous):

Woohoo Im getting it haha

OpenStudy (austinl):

So, what would we have if we combine like terms? :D

OpenStudy (anonymous):

Would it be 62.10= 23+ 1.25?

OpenStudy (austinl):

Nope, it is .5x+.075x right?

OpenStudy (anonymous):

Oops haha

OpenStudy (anonymous):

Im sorry

OpenStudy (austinl):

No need to be sorry :) Happens to the best of us, super simple mistake to make! :D So what would the equation look like now?

OpenStudy (anonymous):

Wait what about the 20 and 3?

OpenStudy (austinl):

20+3=?

OpenStudy (anonymous):

23 So the problem would look like this 62.10= 23+.5x+.75x?

OpenStudy (austinl):

62.10=23+.575x

OpenStudy (anonymous):

Thx for the medal :) How do I give u one?

OpenStudy (anonymous):

Oh okay

OpenStudy (austinl):

You just click best response by a reply of mine, blue button.

OpenStudy (anonymous):

There u go <3

OpenStudy (austinl):

:D thank you ♥

OpenStudy (anonymous):

Okay so what do we do next in this problem?

OpenStudy (austinl):

We would need to get constants on one side, and terms with variables on the other. How would we do this?

OpenStudy (anonymous):

Would we subtract 23 from 62.10?

OpenStudy (austinl):

Indeed we would!

OpenStudy (anonymous):

Finally I get it right haha So the answer after subtracting it would be 39.1?

OpenStudy (austinl):

Yep, we would now have, 39.7=0.575x What would we do to get x by itself?

OpenStudy (anonymous):

We would divide 39.1 by 0.575?

OpenStudy (austinl):

Yep, what do you get?

OpenStudy (anonymous):

68?

OpenStudy (austinl):

There we have it!! :D

OpenStudy (anonymous):

So x=68?

OpenStudy (austinl):

Yep, she can buy 68 pages.

OpenStudy (anonymous):

I luv you thank you so much :)

OpenStudy (austinl):

No problemo, loves ya too :P

OpenStudy (anonymous):

I think I can do the other problems now So Ima go :) Have a good day and kik me back sometime XD

OpenStudy (austinl):

When I am free to kik, your phone or whaterver must always be off. It is always stuck on sending :P Are you not able during the day?

OpenStudy (anonymous):

Nah I share a phone with my brother so I don't get it till nightime

OpenStudy (austinl):

Do you have an iPod?